OpenStudy (anonymous):

Simplify 2 square root 162 with an index of 3?

OpenStudy (anonymous):

did you mean cube root?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[2\sqrt[3]{162}?\]

OpenStudy (anonymous):

YES! sorry i wasnt sure how to word it

OpenStudy (anonymous):

2*5.45=10.90

OpenStudy (anonymous):

\[\sqrt[3]{162}=5.45\]

OpenStudy (anonymous):

My answer choices arent decimals

OpenStudy (anonymous):

162 = 2 * 3*3*3*3=6*3^3 so \[\sqrt[3]{162}=3\sqrt[3]{6}=> 2\sqrt[3]{162}= 2*3\sqrt[3]{6}=6\sqrt[3]{6}\]

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

You're welcome :) Just break down the large number

OpenStudy (anonymous):

what was the tool that you used to show radicals?

OpenStudy (anonymous):

You saw Equation Draw Attach File under the reply box? Use the Equation, they have all symbols you need